# Predictive Reliability Assessment of Distribution Grids with Residential Distributed Energy Resources

### v1

@@ Abstract
Distribution system end users are transforming from passive to active participants, marked by the push towards widespread adoption of edge-level Distributed Energy Resources (DERs). This paper addresses the challenges in distribution system planning arising from these dynamic changes. We introduce a bottom-up probabilistic approach that integrates these edge-level DERs into the reliability evaluation process. Our methodology leverages joint probability distributions to characterize and model the penetration of rooftop photovoltaic (PV) systems and energy storage across a distribution network at the individual residential level. Employing a scenario-based approach, we showcase the application of our probabilistic method using a Monte Carlo Simulation process to assess average system reliability indices and their variations at the user level. To validate our approach, we applied this methodology to the RBTS test system across various adoption scenarios, effectively showcasing the capability of our proposed method in quantifying the variation in end-user reliability indices for each scenario within the distribution system.
@@
Distributed energy resources, predictive reliability assessment, distribution grids.
@@ I Introduction
Distribution systems have transitioned from unidirectional, static, radial systems [1] to bi-directional, dynamic, and meshed systems and evolved to become complex interconnected systems. The importance of distribution system reliability for utilities/distribution system operators (DSO) cannot be overstated, encompassing diverse areas from community welfare to asset management. Beyond their primary function of transporting electricity, distribution systems are foundational in upholding public health and fostering a community’s economic vitality [2]. Consequently, maintaining their efficient performance is not a mere operational necessity but is crucial for the broader well-being of communities. Utilities increasingly apply asset management techniques to their distribution systems to identify factors influencing infrastructure reliability to make informed decisions about their systems’ maintenance, upgrades, and investments [3].
Customer/end user satisfaction, an integral measure for utilities, is directly influenced by system reliability. Utilities ensure a consistent power supply by assessing reliability and fostering an environment conducive to economic development and heightened customer satisfaction. Furthermore, reliability assessments help utilities optimize investments in distribution systems, minimizing the costs of interruptions to both the utility and its customers, leading to a more cost-effective and efficient distribution system [4].
There are two types of reliability assessments: historical and predictive. The objective of historical assessment is to measure and report the system’s actual performance, identifying areas of improvement based on past and current data. On the other hand, predictive reliability assessment aims to quantify potential reliability issues based on projections, scenarios, and anticipated changes in the system. Its objective is to guide the expansion and development of the distribution system in planning and designing systems that can adapt to future changes and challenges. This paper is motivated by the need for a predictive assessment method that integrates the end user DER into the reliability assessment.
@@ I-A Related Works
As DERs became an emerging technology in distribution systems, the focus was on integrating them at the system/feeder level. In this context of reliability assessment, the research focused on integrating these resources into the assessment framework [5]. Both analytical [6, 7, 8, 9, 10] and simulation methods [11, 12, 13, 14, 15, 16, 17], were developed to align with the existing literature. For instance, [6] presents a generalized analytical formulation for reliability evaluation with DGs and microgrids, and [7] proposes an encoded Markov set algorithm for reliability evaluation with both conventional and renewable DG. [8] proposes an analytical method considering the protection strategy, islanding operation, and supply restoration process in the presence of DGs.
Similarly, sequential [11, 12] and non-sequential simulation [15] methods were developed to integrate Distributed energy resources (DERs) into reliability evaluation, considering various aspects such as energy storage integration [11], the islanding dynamics of microgrids [15], and electric vehicle integration [13, 16, 17]. All these approaches evaluated the primary load point reliability indices, namely failure rate ($\lambda$), repair time ($r$), and unavailability ($U$) in the presence of DGs; these load point indices were used to calculate the system reliability indices using the formulation in [18]. [19], present a three level hierarchical reliability assessment framework to integrate DERs both at the system/feeder level and the end user level.
Limitations of the current reliability assessment methods: Behind-the-meter (BTM) DERs allow end users to play a dynamic role as a sink and source based on the operating conditions of the distribution system. Given this new paradigm in the distribution system, using primary load point indices to evaluate reliability is myopic. The following example elucidates the limitations of the current approach. Consider a system with two load points and five residential customers connected to each of them. Load point data is provided in Table I.
@@
Load points customers $\lambda$ (f/y) U (hrs)
@@
The reliability indices for the example system evaluated by the current approach of using the basic load point indices are as follows:
$\text{SAIFI}_{\text{P}}=\frac{\sum\lambda_{i}C_{i}}{\sum C_{i}},\quad\text{SAIDI}_{\text{P}}=\frac{\sum U_{i}C_{i}}{\sum C_{i}}$ (1)
The subscript ’P’ here means ’perceived’ indicating that the reliability metrics evaluated are as seen by the DSO/utility side. The actual customer experience can be quantified using the same reliability indices as follows:
$\text{SAIFI}_{\text{E}}=\frac{\sum_{C}\text{Interruptions experienced }}{\sum C}$ (2)
$\text{SAIDI}_{\text{E}}=\frac{\sum_{C}\text{Interruption durationexperienced }}{\sum C}$ (3)
The subscript ’E’ means experienced indicating that the reliability metrics evaluated are as seen from the customer side of the distribution system. Consider the following two cases for the example system:
Figure 1: Case1: NO DER, Case2: DER at 4,5,9,10.
Case 1: There are no BTM DERs in this case. The failure rate and the unavailability of the load point are equal to the interruption frequency and the restoration time of the end user, respectively. Therefore, the system indices evaluated using the load point approach and those evaluated from customer experience yield the same numerical result. Similar results indicate that DSO’s perceived reliability is same as the reliability experienced by the customers, as shown in Fig. 1.
Case 2: 40% of the customers have BTM DER (customers 4,5,9 and 10). As shown in Fig. 1, for customers 4,5,9 and 10, BTM DER can meet the system demand whenever there is a load point failure, therefore they never experience an interruption. Given this context, the system indices evaluated using the load point approach by the DSO and that evaluated from customer experience yield different numerical results. The disparity in the indices indicates that the reliability perceived by the DSO is not the same as the reliability experienced by the customers. While the example assumes simplistic scenarios, it shows that the current approach of using the basic load point indices to evaluate system reliability does not correctly quantify the reliability value provided by BTM DER assets.
I-B Motivation and Scope of this Work
In the evolving landscape of modern distribution systems, the proliferation of BTM DERs challenges traditional notions of reliability evaluation. Modern energy consumers, driven by rapidly shifting paradigms, demand more than just power supply. Their expectations encompass flexibility, resilience, and active participation in the energy market through DERs. This active involvement can significantly influence customer satisfaction levels, bridging or widening the gap between utility offerings and end-user expectations. Relying on traditional financial evaluations might lead utilities down paths that, while appearing cost-effective at first glance, do not account for the full spectrum of value offered by DERs, potentially resulting in subpar investment decisions. The dynamic integration of DERs affects wear and tear, performance, and asset management strategies. By overlooking their influence, utilities may inadvertently miss signals crucial for timely maintenance, putting the system’s integrity and reliability at risk.
Furthermore, the contemporary energy transition, driven in part by the integration of DERs, offers an unprecedented opportunity to address longstanding disparities in energy access and costs. Disregarding these resources risks perpetuating past inequities, a misstep starkly contrasting the ideals of a modern, inclusive grid. As the role of end-users undergoes a transformation from passive consumption to active production and integration, reliability assessments must capture this significant shift.
This paper builds on the existing literature on distribution system reliability assessment. The proposed method represents a modular addition to the existing assessment methods, designed to integrate seamlessly and augment their functionality. In this bottom-up approach, we first provide a residential system model that can be configured based on the BTM DER. We provide the formulations for the reliability evaluations for the residential system. We developed the probabilistic approach and the required DER penetration modeling techniques for predictive reliability assessment. We applied the proposed method to a modified RBTS test system and critically analyzed the results. This analysis demonstrated the practical application of the approach, revealing its effectiveness in various scenarios. The results underscore the utility of the probabilistic method in enhancing the accuracy and comprehensiveness of reliability assessments in modern distribution systems. They also highlight the method’s ability to account for the complexities and variabilities introduced by DERs, showcasing its potential as a valuable tool for future-oriented energy system planning and management. The contributions of this paper are summarized as follows:
  • We propose a bottom-up probabilistic approach to evaluate the reliability of distribution systems using a general-purpose modular residential system configuration.
  • We define a novel way of modeling DER penetration utilizing probability distributions to model micro-level penetration at the granularity of the residential level.
  • We propose a Monte Carlo Simulation method to estimate distribution system reliability indices and their underlying distributions at the residential level.
@@ I-B Motivation and Scope of this Work
The remainder of the manuscript is organized as follows: Section 2 describes the residential system reliability evaluation. Section 3 proposes the probabilistic approach to system reliability evaluation. Section 4 introduces the DER penetration model. Section 5 describes the simulation method to evaluate system reliability and application of the methodology to RBTS test system.
@@ II-A2 ES system model
Energy storage is modeled using the State of Charge (SOC) of the system where SOC is percentage that represents the fraction of the maximum capacity.
@@ III Probabilistic approach to system reliability assessment
To address the disparity in load magnitude, PV system sizes, and energy storage capacities among different residential systems, we normalize these variables with the peak load ($l_{c,p}$), yielding:
@@ IV-A Macro vs Micro Perspectives
These macro penetration metrics $\phi_{\text{PV}}$, $\phi_{\text{ES}}$ provide a holistic view of how saturated a given system or region is with the adopted DER technology. While useful, this high-level perspective can obscure variations and patterns at the residential levels. In this rapidly evolving landscape of edge-level DERs, understanding the penetration of DER can no longer remain a macro affair. Instead, capturing the intricacies of individual behaviors, choices, and motivations is paramount as the actual dynamics of energy transition play out at this granular edge level. By modeling the penetration at the micro-level, one can peel back the layers of aggregate statistics to reveal the vibrant mosaic of individual decisions and factors that drive the bigger picture.
@@ IV-B Quantitative Modeling of Micro-level Penetration
V Application of the probabilistic approach
V-A MCS Simulation
We propose a Monte Carlo simulation (MCS) framework that leverages the probabilistic model discussed earlier to estimate the distribution of reliability indices such as AIF and AID across the distribution system. The following steps outline the MCS:
  1. 1.
    Sample Generation: Draw samples from the joint probability distribution of X and Y.
  2. 2.
    Index calculation: Each sample corresponds to a specific configuration of X and Y for the residential system. Evaluate the desired reliability metric for every unique system configuration derived from the samples.
  3. 3.
    Stopping criterion: Employ the confidence interval width as the stopping criterion. Continue simulations until the width of the CI is within a predefined precision threshold.
The detailed simulation flowchart is shown in Fig. 3. The steps described above are the generalized MCS steps for the abstraction where the reliability characteristics of the load points are uniform across the distribution system. In the case of more granular abstraction layers, such as the one where each load point has different reliability characteristics, the same set of steps detailed earlier are used to evaluate the index at each load point, and the system index is then calculated using the weighted average of the load point indices as expressed in (33).
Figure 3: MCS flowchart.
V-B Marginal Distributions for X and Y
Field penetration data provides a solid foundation for accurately modeling DER penetration. However, a significant challenge arises due to the shallow adoption of BTM DER, especially in energy storage in the current energy landscape. This shallow adoption hinders our ability to construct meaningful, data-driven adoption scenarios. Consequently, we rely heavily on hypothetical constructions and theoretical models. By shaping distributions with these theoretical methods, we can approximate the specific patterns that might emerge in reality. These hypothetical models are tools for current understanding and are invaluable in predicting future Photovoltaic (PV) penetration scenarios. By integrating forecasts and considering potential shifts in policy and technology, we can use these theoretical constructs to anticipate and prepare for changes in the energy landscape despite the current limitations in real-world data. It is worth noting, however, that the core objective of this paper diverges slightly. Rather than focusing on the details of creating these distributions, our primary interest lies in developing a framework to utilize these distributions to derive scenario-based insights on system reliability.
@@ V-B Marginal Distributions for X and Y
Similarly, with residential energy storage, the popular residential energy storage solutions have capacities of 13.5kWh [21]. Considering this system as an average, it can support the peak load of an average household for approximately 3.375 hours. By extending the range from 0 to 6.75 hours, we encompass the full breadth of storage capacities, considering that some homes may opt not to have energy storage. In contrast, others may install systems that offer double the average duration of peak load coverage, thus ensuring preparedness for scenarios from short-term outages to extended periods without grid power.
These ranges are not arbitrary but thoughtfully chosen to represent the variability inherent in residential energy solutions. The average values give us a baseline, while the extended ranges ensure we capture the real-world fluctuations—both in terms of solar generation potential and energy storage capacity—thus providing a more comprehensive picture that homeowners can relate to and plan for.
@@
where $\alpha=0.75$ and $\beta=5$. The distribution is skewed towards the lower end of the range and the support is shifted from the standard [0,1] interval to [0, 3.5] for X and [0, 6.75] for Y by the $loc$ and $scale$ parameters.
@@
  • Y:Energy storage systems are not widely adopted. Among those with them, the storage capacity is minimal, possibly due to high costs, lack of awareness, or limited perceived need.
@@
  • Y: Many homes have adopted energy storage systems, but their installed capacities vary greatly. While some homes might have minimal storage, others might approach the upper end of the capacity spectrum.
@@
  • Y: Most homes with energy storage systems have capacities close to the median value. This could be due to specific popular product offerings or household energy storage needs.
@@
  • Y: In this scenario, homes that adopt energy storage systems tend to install capacities on the higher end of the spectrum. Reasons could include larger homes, areas with unreliable grid power, or incentives that make larger systems more attractive.
V-C Joint adoption scenarios

### v2

@@ Abstract
The growing adoption of behind-the-meter distributed energy resources (DERs) such as rooftop photovoltaic (PV) systems and energy storage (ES) is transforming end users from passive consumers to active participants in distribution systems, creating new challenges for reliability evaluation. This paper proposes a bottom-up probabilistic framework that integrates residential DER adoption into predictive reliability assessment. The framework models PV and ES adoption using joint probability distributions at the customer level, incorporates component reliability, and employs an adaptive Monte Carlo simulation to estimate not only mean values of reliability indices but also their distributional variability across customers. The approach is demonstrated on the RBTS Bus 4 system under sixteen joint adoption scenarios. Results show that high PV–storage adoption can reduce system-average SAIFI and SAIDI by more than 90% compared to the baseline, while shallow adoption may worsen interruption frequency for some users. These findings highlight the importance of capturing adoption heterogeneity and distributional outcomes in reliability assessment, providing utilities and regulators with a robust tool for planning in DER-rich distribution grids.
@@
Distributed energy resources, predictive reliability assessment, distribution systems.
@@ I Introduction
Distribution systems have transitioned from unidirectional, static, radial systems [1] to bi-directional, dynamic, and meshed systems, evolving into complex interconnected networks. The importance of distribution system reliability for utilities and distribution system operators (DSOs) cannot be overstated, encompassing areas from community welfare to public health and economic vitality [2]. Utilities increasingly apply asset management techniques to their distribution systems to identify factors influencing reliability and to make informed decisions about maintenance, upgrades, and investments [3].
Customer satisfaction, an integral measure for utilities, is directly influenced by system reliability. Utilities ensure a consistent power supply by assessing reliability and fostering an environment conducive to economic development and heightened customer satisfaction. Reliability assessments also help utilities optimize infrastructure investments, minimizing the costs of interruptions to both utilities and customers, and thereby contributing to a more cost-effective and efficient distribution system [4].
Two forms of reliability assessment are commonly applied. Historical assessment evaluates realized system performance using past and current data, while predictive assessment quantifies potential reliability outcomes under projected changes, scenarios, or future contingencies. Predictive assessment is particularly critical for planning, as it enables system designs that can adapt to emerging challenges. This paper is motivated by the need for predictive reliability methods that explicitly integrate end-user distributed energy resources (DERs), which are reshaping the nature of distribution system performance.
@@ I-A Related Works
Analytical approaches. Early research on distribution system reliability with DERs emphasized analytical formulations. Examples include generalized reliability evaluation with distributed generation and microgrids [5], encoded Markov set algorithms for hybrid renewable–conventional DG systems [6], and methods explicitly considering protection strategy, islanding, and restoration in the presence of DG [7]. These methods provide closed-form insights and computational efficiency but are typically limited by state-space growth and difficulty in capturing the stochastic dynamics of intermittent DERs at scale. More recent analytical studies [8, 9, 10] have extended recursive and state enumeration formulations to active distribution networks, microgrids, and DER uncertainty. While these models improve tractability for certain contexts, they continue to focus on mean load-point or system indices and cannot capture heterogeneity introduced by diverse end-user adoption.
Simulation approaches. Monte Carlo simulation frameworks provide a more flexible alternative, enabling sequential and non-sequential representations of DER operation. Sequential simulation has been used to evaluate energy storage adequacy and microgrid islanding [11, 12], while non-sequential methods incorporate DER and EV adoption under diverse outage scenarios [13, 14, 15]. More recent works [16, 17, 18] have applied MCS to customer reliability, PHEV integration, and demand response. However, these studies largely report average indices such as SAIFI and SAIDI at the feeder or system level, and seldom explore the distributional variability across customers or include component reliability of PV and storage devices.
In summary, analytical approaches offer compact formulations but suffer from state-space limitations and difficulty in modeling stochastic DER adoption, while simulation approaches provide flexibility but are computationally demanding and typically report only average reliability indices. Neither category fully captures the distributional impacts of heterogeneous end-user DER adoption or integrates component-level PV and storage reliability, which motivates the need for a new framework developed in this work.
I-B Motivation and Scope of this Work
Motivating example: [DSO vs. customer reliability] Behind-the-meter (BTM) DERs allow end users to act as both sinks and sources of power depending on operating conditions. This fundamentally changes how reliability should be measured. Consider a simple distribution system with two load points, each serving five residential customers, as summarized in Table I.
@@ I-B Motivation and Scope of this Work
Load point Customers $\lambda$ (f/yr) U (h/yr)
@@
Using the traditional approach, reliability indices are evaluated with the basic load point data as
$\text{SAIFI}_{P}=\frac{\sum\lambda_{i}C_{i}}{\sum C_{i}},\quad\text{SAIDI}_{P}=\frac{\sum U_{i}C_{i}}{\sum C_{i}},$ (1)
where the subscript “P” denotes perceived indices as reported by the DSO/utility. The actual customer experience can be represented by
$\text{SAIFI}_{E}=\frac{\sum_{C}\text{Interruptions experienced}}{\sum C},$ (2)
$\text{SAIDI}_{E}=\frac{\sum_{C}\text{Interruption duration experienced}}{\sum C},$ (3)
where the subscript “E” denotes experienced indices at the end-user level.
Figure 1: Case 1: no DER; Case 2: DER at {4,5,9,10}.
Case 1: Without DERs, perceived and experienced reliability coincide; each customer is equally exposed to interruptions.
Case 2: When 40% of customers adopt BTM DERs (customers 4, 5, 9, and 10), those customers no longer experience interruptions. As shown in Fig. 1, the indices calculated from the DSO’s perspective differ significantly from those experienced by customers. This example highlights that load-point–based methods cannot quantify the reliability value provided by BTM DERs and motivates the need for an approach that captures customer-level outcomes.
Motivation: As residential DER adoption accelerates, reliability assessment must move beyond average indices reported from the utility perspective. Modern customers expect not only continuity of supply but also resilience, flexibility, and recognition of their DER contributions. Ignoring heterogeneity in adoption risks undervaluing customer investments and misguiding planning decisions. Predictive methods must therefore integrate probabilistic adoption models and account for PV and ES component reliability, enabling fair and forward-looking reliability evaluation.
Scope of this work. This paper proposes a modular, bottom-up reliability assessment framework that augments conventional methods. The framework models residential systems with configurable BTM DER adoption, integrates PV and ES component reliability, and applies adaptive Monte Carlo simulation to estimate not only mean indices but also the distribution of customer-level outcomes. The approach is demonstrated on a modified RBTS Bus 4 system with 16 joint PV–ES adoption scenarios.
Contributions. The contributions of this paper are summarized as follows:
  • We propose a bottom-up probabilistic framework that evaluates distribution system reliability by aggregating residential-level DER adoption into system indices, in contrast with conventional top-down approaches that compute system indices directly from load-point averages reported at the utility level.
  • We introduce a novel adoption-modeling approach that uses probability distributions to represent micro-level DER penetration.
  • We incorporate PV and ES component reliability into system-level reliability indices.
  • We employ an adaptive Monte Carlo simulation method to estimate not only SAIFI/SAIDI means but also their distributional variability across customers.
  • We demonstrate the framework on a modified RBTS Bus 4 case study, analyzing 16 joint PV–ES adoption scenarios.
@@
The remainder of this paper is organized as follows: Section 2 describes the residential system reliability evaluation. Section 3 presents the probabilistic formulation. Section 4 introduces the DER penetration models. Section 5 details the Monte Carlo simulation method and its application to the RBTS system.
@@ II-A2 ES system model
ES is modeled using the State of Charge (SOC) of the system where SOC is percentage that represents the fraction of the maximum capacity.
@@ III Probabilistic approach to system reliability assessment
Here $g(\cdot)$ and $h(\cdot)$ denote the reliability mapping functions introduced in the Monte Carlo formulation. The function $g(\cdot)$ maps a realization of adoption ratios $(X,Y)$ and outage conditions to the interruption frequency contribution (AIF) for a representative customer, i.e., the number of interruptions experienced during that trial. Similarly, $h(\cdot)$ maps the same inputs to the interruption duration contribution (AID), i.e., the number of outage hours incurred by the customer.
@@
To address the disparity in load magnitude, PV system sizes, and ES capacities among different residential systems, we normalize these variables with the peak load ($l_{c,p}$), yielding:
@@ IV-A Macro vs Micro Perspectives
These macro penetration metrics $\phi_{\text{PV}}$, $\phi_{\text{ES}}$ provide a holistic view of how saturated a given system or region is with the adopted DER technology. Although useful, this high-level perspective can obscure variations and patterns at the residential levels. In this rapidly evolving landscape of edge-level DERs, understanding the penetration of DER can no longer remain a macro-problem. Instead, capturing the intricacies of individual behaviors, choices, and motivations is paramount as the actual dynamics of the energy transition plays out at this granular edge level. By modeling the penetration at the microlevel, one can peel back the layers of aggregate statistics to reveal the vibrant mosaic of individual decisions and factors that drive the bigger picture.
@@ IV-B Quantitative Modeling of Micro-level Penetration
IV-C Marginal Distributions for X and Y
Field penetration data provides a solid foundation for accurately modeling DER penetration. However, a significant challenge arises due to the shallow adoption of BTM DER, especially in ES in the current energy landscape. This shallow adoption hinders our ability to construct meaningful, data-driven adoption scenarios. Consequently, we rely heavily on hypothetical constructions and theoretical models. By shaping distributions with these theoretical methods, we can approximate the specific patterns that might emerge in reality. These hypothetical models are tools for current understanding and are invaluable in predicting future Photovoltaic (PV) penetration scenarios. By integrating forecasts and considering potential shifts in policy and technology, we can use these theoretical constructs to anticipate and prepare for changes in the energy landscape despite the current limitations in real-world data. It is worth noting, however, that the core objective of this paper diverges slightly. Rather than focusing on the details of creating these distributions, our primary interest lies in developing a framework to utilize these distributions to derive scenario-based insights on system reliability.
@@ IV-C Marginal Distributions for X and Y
Similarly, with residential ES, the popular residential ES solutions have capacities of 13.5kWh [20]. Considering this system as an average, it can support the peak load of an average household for approximately 3.375 hours. By extending the range from 0 to 6.75 hours, we encompass the full breadth of storage capacities, considering that some homes may opt not to have ES. In contrast, others may install systems that offer double the average duration of peak load coverage, thus ensuring preparedness for scenarios from short-term outages to extended periods without grid power.
These ranges are not arbitrary but thoughtfully chosen to represent the variability inherent in residential energy solutions. The average values give us a baseline, while the extended ranges ensure we capture the real-world fluctuations, both in terms of solar generation potential and energy storage capacity, thus providing a more comprehensive picture that homeowners can relate to and plan for.
@@
where $\alpha=0.75$ and $\beta=5$. The distribution is skewed towards the lower end of the range and the support is shifted from the standard [0,1] interval to [0, 3.5] for X and [0, 6.75] for Y by the locloc and scalescale parameters.
@@
  • Y:ES systems are not widely adopted. Among those with them, the storage capacity is minimal, possibly due to high costs, lack of awareness, or limited perceived need.
@@
  • Y: Many homes have adopted ES systems, but their installed capacities vary greatly. While some homes might have minimal storage, others might approach the upper end of the capacity spectrum.
@@
  • Y: Most homes with ES systems have capacities close to the median value. This could be due to specific popular product offerings or household ES needs.
@@
  • Y: In this scenario, homes that adopt ES systems tend to install capacities on the higher end of the spectrum. Reasons could include larger homes, areas with unreliable grid power, or incentives that make larger systems more attractive.
IV-D Joint adoption scenarios